OpenStudy (anonymous):

The solution to a system of three equations in three variables is ________ one point. A. always B. sometimes C. never

OpenStudy (zzr0ck3r):

what do you think?

OpenStudy (zzr0ck3r):

how many solutions to x+y+z=3 x+y+z=3 x+y+z=3 how many solutions to x+0y+0z=3 0x+0y+z=4 x+y+z=17

OpenStudy (anonymous):

sometime i think

OpenStudy (zzr0ck3r):

correct